# SIPDO: Closed-Loop Prompt Optimization via Synthetic Data Feedback

### v2

@@ 4 Experiments
We test SIPDO on four main datasets to measure its resilience across different domains and reasoning tasks. We include all 4689 instances from six BIG-Bench tasks(Tables of Penguins, Geometric Shapes, Epistemic Reasoning, Object Counting, Temporal Sequences, and Causal JudgmentSrivastava et al. (2022)). To assess logical reasoning, we sample 600 examples from the depth-5 subset of ProofWriter with a balanced label distribution Tafjord et al. (2021), use 204 test examples from FOLIO that require first-order inference over short passages Han et al. (2024), and select the 500 most challenging 5-hop scenarios from the fictional-character version of PrOntoQA Saparov and He (2022).
@@ 4.1 EXPERIMENTAL SETUP
4.1.1 Baselines
We compare SIPDO with four existing prompt optimization strategies:
Chain of Thought (CoT) Suzgun et al. (2022) improves LLM reasoning by explicitly guiding models through step-by-step decomposition of complex tasks.
Automatic Prompt Engineer(APE) Wang et al. (2023) automatically generates and refines prompts using a Monte Carlo search based on model feedback, improving instruction quality with minimal human intervention.
PromptAgent Zhou et al. (2022b) formulates prompt optimization as a strategic planning task, using Monte Carlo Tree Search (MCTS) to explore the prompt space. It refines prompts based on model errors and feedback, aiming to generate prompts with expert-level quality through systematic trial-and-error refinement.
Neuro-Symbolic Pan et al. (2023) combines neural networks with symbolic rule-based reasoning by transforming LLM outputs into structured symbolic representations. It enables models to handle complex logical inference tasks by integrating natural language understanding with robust symbolic processing, offering a more versatile and human-like approach.
REVOLVE Zhang et al. (2024) tracks the way an LLM’s responses evolve over successive iterations and updates the prompt with this trajectory-aware signal, giving the optimizer a second-order-style view of progress. By looking beyond the most recent feedback, it avoids local optima and usually reaches a higher-quality prompt in fewer steps.
4.2 Results and analyze
4.2.1 Synthetic Data Generation
We fix the difficulty budget at $c=10$ for all benchmarks except Penguins and Geometry from BIG-Bench, where we double it to $c=20$ to accommodate the intrinsically harder structure of these tasks. For every dataset, we also tie the number of training iterations to the difficulty level by setting $t=c$; the optimiser therefore sees one harder tier at each pass. A sample of synthetic Causal-Judgement data produced by SIPDO at difficulty level 8(on a 1-10 scale) is shown in the text box below with additional examples in Appendix  D. These data ensure a highly consistent format with a well-organized structure and clear hierarchical distinctions. Each output adheres strictly to predefined standards, guaranteeing data stability and integrity while significantly reducing the cost of subsequent cleaning and adjustments. Additionally, the structured format enhances readability and parsability, streamlining analysis and processing while improving data reusability and reliability.
Generated Question: In a small town, there is a bakery that makes the best pastries. Every morning, the bakery opens at 7:00 am, and a line of customers forms outside. The bakery owner has a rule that only one person can enter at a time to maintain order. One day, two customers, Alice and Bob, arrive at the same time. Alice follows the rule and waits outside, but Bob ignores the rule and enters the bakery while another customer is still inside. The bakery becomes overcrowded, and a shelf of pastries falls over, ruining the day’s batch. Did Bob cause the pastries to be ruined?
Generated Answer: Yes
By generating complex and challenging scenarios gradually, the data stresses the current prompt and express fresh failure modes, improving its ability to handle a wide range of difficulties effectively. This ensures that the prompt becomes more adaptable and robust as it reaches the hardest tiers. These results strongly validate the effectiveness of structured output in standardizing data generation and improving overall data quality, while fostering continuous advancements in prompt optimization.
4.2.2 Implementation
Geometry data generation and prompt improvement.
@@ 4.2.2 Implementation
Constructing complex and irregular shapes, including irregular shapes and polygons, cannot be effectively achieved using few-shot methods, so we introduce three measures specifically for SVG(Scalable Vector Graphics) path generation in geometry task: (1) precision normalization: every coordinate produced by the generator is rounded to two decimal places, eliminating the floating-point drift that causes downstream parsers to miscount line (L) and arc (A) instructions. (2) template-guided retrieval: before generation, a retriever selects a canonical path template whose instruction pattern matches the target shape (e.g., “4 L” for a rectangle, “1 A” for a sector). The generator then perturbs only the vertex coordinates, guaranteeing syntactic correctness while still exposing the prompt to unseen geometries. (3) reverse-generation check: because the label (shape name) is known a prior, we parse the produced SVG with a deterministic rule-based decoder that counts L and A commands; if the inferred label disagrees with the intended one, the sample is rejected and regenerated. The generated SVG path example is shown below. For further prompt refinement, we provide prompt templates in Appendix C.
Generated SVG Path: <path d="M 23.45,45.78 L 78.32,45.78 L 78.32,78.56 L 23.45,78.56 L 23.45,45.78"/>
Generated Answer:
"target_scores": {
"rectangle": 1, "sector": 0, "triangle": 0, "circle": 0, "heptagon": 0, "hexagon": 0, "kite": 0, "line": 0, "octagon": 0, "pentagon": 0 }
4.2.3 Comparison with prompting baselines
We tested SIPDO by different LLMs such as GPT-4o, GPT-4o-mini, Gemini-1.5-flash, and Gemini-1.5-pro on different datasets, including BIG-Bench, FOLIO, PrOntoQA, and ProofWriter.
Table 1: Results on BIG-Bench tasks across multiple LLMs. SIPDO consistently outperforms standard prompting baselines (CoT, APE, PromptAgent) across most tasks and models, demonstrating generalization and effectiveness of the optimization.
Model Method Accuracy (%) Avg. (Comparative Acc.)
Penguins Geometry Epistemic Obj. Count Temporal Causal
@@ 4.2.3 Comparison with prompting baselines
BIG-Bench. We evaluated SIPDO on six BIG-Bench tasks. As shown in Table 1, GPT-4o and GPT-4o-mini demonstrate particularly strong performance in Temporal Reasoning, Object Counting, and Causal Judgment. While Geometry exhibits comparable accuracy across GPT-4o and GPT-4o-mini, SIPDO achieves the highest overall accuracy for GPT-4o-mini, Gemini-1.5-flash, and Gemini-1.5-pro. For GPT-4o, SIPDO performs at a near-best level, trailing PromptAgent by only 0.01, yet still demonstrating that LLMs benefit from synthetic data generation for reasoning improvements, whereas other methods primarily rely on existing datasets. These results further highlight the advantages of LLM-driven data augmentation in enhancing logical, numerical, and causal reasoning capabilities.
FOLIO, PrOntoQA, and ProofWriter. We evaluated SIPDO, REVOLVE, CoT, Neuro-Symbolic, and a plain baseline prompt on FOLIO, PrOntoQA, and ProofWriter with both GPT-4o and GPT-4o-mini, assessing the methods’ ability to perform structured logical reasoning. As shown in Table 2, SIPDO achieves the highest average accuracy score and outperforms all approaches on FOLIO and PrOntoQA. In PrOntoQA, SIPDO surpasses all methods, demonstrating its capability to generate structured logical proofs. Similarly, for FOLIO, SIPDO outperforms Neuro-Symbolic, CoT, and REVOLVE, further validating its effectiveness in formal logic inference.
Table 2: Results on ProofWriter, FOLIO, and PrOntoQA by Neuro-Symbolic, CoT, REVOLVE, SIPDO, and Baseline Prompting methods across GPT-4o and GPT-4o-mini.
@@
Tasks Baseline Neuro-S CoT REVOLVE SIPDO Baseline Neuro-S CoT REVOLVE SIPDO
@@
FOLIO 71.2 79.2 72.6 65.7 83.2 51.2 73.2 69.3 62.8 81.1
PrOntoQA 80.4 85.2 95.6 85.4 96.3 74.6 79.3 89.3 83.4 91.3
Average 70.0 82.0 80.2 68.4 86.4 59.5 77.4 73.5 64.9 83.9
While neuro-symbolic reasoning remains the best performer on ProofWriter, SIPDO achieves highly competitive results on ProofWriter, trailing by only 0.004 on GPT-4o-mini and 0.02 on GPT-4o, underscoring its strong adaptability to structured reasoning tasks. Crucially, unlike neuro-symbolic approaches that rely on predefined rule-based datasets, SIPDO is trained entirely on generated synthetic data, demonstrating the effectiveness of LLM-driven data augmentation for enhancing logical inference across diverse reasoning benchmarks.
4.2.4 Prompt generalization
The whole process of the prompt improvement on the Penguin data is shown in Appendix  C. Follow the three modules-error analysis, recommendation, and refinement-and outputs a revised prompt. The prompts generated by SIPDO outperform other baselines, validating the effectiveness of SIPDO in optimizing prompt design and enhancing overall performance. By employing a LEAST-TO-MOSTZhou et al. (2022a) approach in prompt generation, the LLM to reason in a structured and incremental manner. After each iteration, the prompts are evaluated using real datasets. If the accuracy falls below a predefined threshold, the prompt is refined based on incorrectly answered questions to enhance its effectiveness. All prompts can be viewed in Appendix  A
4.3 Ablation Study
Model Penguins Geometry Epistemic Obj.Cnt. Temporal Causal Avg.
@@ 4.3 Ablation Study
Difficulty Gradient.
To assess the contribution of the difficulty gradient, we conduct an ablation study by comparing without difficulty gradient. As Table  3 shows, every BIG-Bench sub-task suffers when the difficulty gradient is absent. On average, GPT-4o loses $17.3\%$ accuracy, while the weaker GPT-4o-mini drops $24.3\%$, confirming that smaller models depends even more on the difficulty gradient. The steepest declines appear on tasks Object Counting (40.9 % and 54.4 %) and Geometric Shapes (17.2 % and 35.1 %). Even comparatively simple tasks—Temporal Sequences and Epistemic Reasoning—still lose up to 6 %. These findings underline that a progressive difficulty gradient is essential: it systematically uncovers a prompt’s blind spots and lets the optimizer repair them before moving on to harder examples. The generated prompts tend to be easier and shorter without difficulty gradient placed, often failing to capture complex reasoning patterns(details in Appendix  B).
One-Shot Extremes.
We experimented with replacing the step-wise difficulty gradient by a one-shot extremes sampler that tells the generator to create the most unusual examples. On our synthetic suites this shortcut delivered no measurable gain. The “extreme” samples were either solved instantly or only slight perturbations of original cases, leaving the optimizer with no fresh errors to exploit. We suspect the idea will pay off in real-world corpora—financial statements, medical notes, legal filings—where genuine edge cases abound and can expose blind spots that our synthetic tasks do not capture.
Reflection Check.
We also inserted a reflection check that re-parses each synthetic pair and discards it if the generator’s answer does not match a rule-based output. Because our generator is already conditioned on the ground-truth label (Assumption A1) and draws heavily on templated examples, most outputs are self-consistent. We, therefore, disable the check in the main pipeline for efficiency, but note that it would be prudent to re-enable it when moving to noisier domains.
@@ 5 Conclusion
In summary, we introduce SIPDO, a method that transforms data augmentation into a real-time feedback signal to enhance prompt optimization. By integrating a data generator that synthesizes progressively more difficult examples with an auto-prompt optimizer that refines prompts, SIPDO systematically identifies and resolves prompts’ weaknesses. This iterative feedback-driven loop improves prompt robustness and performance across diverse reasoning tasks and benchmarks. Empirical evaluations show that SIPDO achieves significant accuracy gains, outperforming several prompt optimization baselines. Additionally, we provide theoretical backing by demonstrating bounded worst-case error once training reaches stability, further supporting SIPDO’s reliability in practice. By enabling LLM systems to autonomously recognize and correct their own shortcomings, SIPDO presents a scalable and efficient strategy toward adaptive, self-improving models capable of generalizing reliably across unseen domains and increasingly complex challenges.
Limitations
Although our multi-agent framework shows clear gains on standard reasoning benchmarks, two practical gaps remain. First, all experiments were run on clean public datasets; we have not yet tested SIPDO on messier, domain-specific collections such as financial filings or clinical notes, where labeling rules and text quality vary widely. Second, the current loop relies on repeated calls, which can be slow and costly at scale. Evaluating on real-world data and exploring lighter generator/critic models are therefore important directions for future work.
Appendix A Optimzed Prompts for different tasks

### v3

@@ 4.1 Experimental Setup
We use following baseline methods are for comparison across the different datasets and benchmarks: Chain of Thought (CoT) (Suzgun et al., 2022) guides models through explicit step-by-step reasoning; Automatic Prompt Engineer (APE) (Wang et al., 2023) refines prompts via Monte Carlo search and model feedback; PromptAgent (Zhou et al., 2022) also uses Monte Carlo Tree Search to iteratively improve prompts; Neuro-Symbolic (Pan et al., 2023) converts LLM outputs into structured forms for rule-based inference; TextGrad (Yuksekgonul et al., 2024) treats textual feedback as a first-order gradient for prompt updates; Momentum-Enhanced TextGrad (Yuksekgonul et al., 2024) adds momentum, enlarging updates when feedback aligns; REVOLVE (Zhang et al., 2024) adjusts prompts using the trajectory of model responses as a second-order-style signal; and ANN (Ma et al., 2025) models agent collaboration as a layered neural network.
We test SIPDO on five main datasets across reasoning tasks:
BIG-Bench. We include all 4689 instances from six BIG-Bench tasks: Penguins In a Table, Geometric Shapes, Epistemic Reasoning, Object Counting, Temporal Sequences, and Causal Judgment (Srivastava et al., 2022). For these tasks, SIPDO is compared to Chain of Thought (CoT) (Suzgun et al., 2022), Automatic Prompt Engineer (APE) (Wang et al., 2023), and PromptAgent (Zhou et al., 2022).
Logical Reasoning Tasks. To assess logical reasoning, we sample 600 examples from the depth-5 subset of ProofWriter with a balanced label distribution (Tafjord et al., 2021), use 204 test examples from FOLIO that require first-order inference over short passages (Han et al., 2024), and select the 500 most challenging 5-hop scenarios from the fictional-character version of PrOntoQA (Saparov & He, 2022). For these tasks, SIPDO is compared to Chain of Thought (CoT) (Suzgun et al., 2022), Neuro-Symbolic (Pan et al., 2023), and REVOLVE (Zhang et al., 2024).
MMLU (Massive Multitask Language Understanding). To test expert-level factual knowledge and problem solving in LLMs, we evaluate on MMLU (Hendrycks et al., 2020), focusing on college-level subject areas: Biology (114 instances), Computer Science (100 instances), and Machine Learning (112 instances). For this benchmark, SIPDO is compared with TextGrad (Yuksekgonul et al., 2024), REVOLVE (Zhang et al., 2024), and ANN (Ma et al., 2025).
4.2 Implementation
Data Generation and Prompt Improvements. We specify a maximum level $c$ and data generated with level of difficulties in prior iterations so that model is aware of the difficulty level of each previous example. To illustrate this, we provide a detailed example of generating a Causal Judgment QA pair from BIG-Bench below.
Causal Judgment Prompt Template [System Input]: You are an expert in generating logical causal-attribute questions and answers. Your task is to generate one pair of causal attributions or causation. One is a causation statement, and another is a non-causation statement. [User Input]: Guidelines: 1. Make sure that the data generated is different. 2. Use clear and direct words in the question, avoiding overly complex phrasing, trickiness, or ambiguity. 3. Do not make the logic in the statement overly complicated; the statement and question should be understood at a fair level. 4. Be creative and diverse. Only follow the logic and flow of the given examples, but be creative and diverse with the content and not limited to the given example. 5. Only one data instance needs to be generated each time. 6. Generate only one output: the generated content should strictly follow the output format from the examples below. 7. Difficulty should increase with each iteration with total difficulty level: {max difficulty level} (current difficulty level: {c}). 8. Make sure the generated data is sufficient and robust without any errors, especially logic. 9. The generated samples are suppose to challenge the model’s ability to reason and answer the question correctly.
Past generated samples: {Generated data with difficulty} Below are the Examples with expected data format:
{True Data 1}
{True Data 2}
We set the difficulty budget at $c=10$ for all benchmarks except Penguins In a Table and Geometric Shapes from BIG-Bench, where $c=25$ accommodates their complex reasoning. The number of training iterations is tied to the difficulty level, with $t=c$ to ensure progressively harder samples. The model temperature is set to $0.5$ for data generation to maintain coherence. We fix the target label $\tilde{y}$ and prompt the model to generate a matching question for data validity. For challenging MMLU benchmarks, three expert agents review each generated item, and only those with unanimous approval are passed to the auto-prompter. This minimizes hallucinations and ensures accurate question–answer pairs. Examples of synthetic BIG-Bench Causal Judgment data at different difficulty levels are shown in Appendix E.
For further prompt optimization, we set the model temperature to 0.0 during error analysis and improvement steps to ensure deterministic, high-quality outputs. However, we use a higher temperature of 0.7 for generating recommendations, which encourages the model to produce more diverse and creative suggestions. This combination stabilizes the error analysis and prompt improvement phases while simultaneously enhancing the breadth and quality of recommendations. The full prompt-improvement process for BIG-Bench Penguins In a Table is in Appendix D.
@@ 4.2 Implementation
Geometric Data Generation. Constructing complex or irregular shapes exceeds the limits of few-shot methods, so we introduce three safeguards for SVG path generation in the geometry task: (1) precision normalization—each coordinate is rounded to two decimal places, preventing floating-point drift that makes downstream parsers miscount line (L) and arc (A) commands; (2) template-guided retrieval—a retriever selects an SVG path template whose instruction pattern matches the target shape (e.g., “4 L” for a rectangle, “1 A” for a sector), and the generator perturbs only the vertex coordinates, ensuring syntactic correctness while adding variety; (3) reverse-generation check—because the shape label is known in advance, a rule-based decoder parses the generated path, tallies L/A commands, and rejects any sample whose inferred label disagrees. Examples appear in Appendix E.
Model Method Accuracy (%) Avg. Acc. (%) (Comparative Acc.)
Penguins Geome. Epistemic Obj. Count Temporal Causal
@@
Table 2: Results on BIG-Bench tasks across multiple LLMs. SIPDO consistently outperforms standard prompting baselines (CoT, APE, PromptAgent) across most tasks and models, demonstrating generalization and effectiveness of the prompt optimization by synthetic data feedback.
4.3 Results and Analysis
We tested SIPDO across various LLMs with temperature of 0.0 on all benchmarks, including BIG-Bench, FOLIO, PrOntoQA, ProofWriter, and three subjects from MMLU. Specifically, we ran SIPDO on GPT-4o, GPT-4o-mini, Gemini-1.5-flash, and Gemini-1.5-pro, so that all synthetic‐data calls and prompt refinements in result were driven by the same model.
MMLU. We first evaluate SIPDO on three subjects from the MMLU benchmark: Machine Learning, College Biology, and College Computer Science. As shown in Table 1, SIPDO achieves the highest results in all three subjects, while TextGrad and REVOLVE tie in the Biology subject. Specifically, we set up the SIPDO pipeline using the GPT-4o model, except for the GPT-4o-mini during the testing phase, in order to provide greater insight into the reasoning weaknesses of LLMs.
BIG-Bench. We then evaluate SIPDO on six BIG-Bench tasks As shown in Table 2, GPT-4o, GPT-4o-mini, and Gemini-1.5-flash demonstrate particularly strong performance in Temporal Reasoning, Object Counting, Penguins In a Table, and Causal Judgment. While Geometric Shapes exhibits comparable accuracy across GPT-4o and GPT-4o-mini, SIPDO achieves the highest overall accuracy across all LLMs except for GPT-4o, trailing PromptAgent by only 0.1%, yet still demonstrating that LLMs benefit from synthetic data generation for reasoning improvements, whereas other methods primarily rely on existing datasets. To further illustrate the generated different difficulty data in Causal Judgment tasks as iteration and difficulty increase, the actual logical turns display a monotonic trend in figure  2.
Figure 2: Generated BIG-Bench Causal Judgment task in different difficulties
FOLIO, PrOntoQA, and ProofWriter. We further test SIPDO on FOLIO, PrOntoQA, and ProofWriter, assessing the methods’ ability to perform structured logical reasoning. As shown in Table 3, SIPDO outperforms all approaches on FOLIO and PrOntoQA and achieves the highest average accuracy. In PrOntoQA, SIPDO surpasses all methods, demonstrating its capability to generate structured logical proofs. Similarly, for FOLIO, SIPDO outperforms Neuro-Symbolic, CoT, and REVOLVE, further validating its effectiveness in formal logic inference.
While neuro-symbolic reasoning remains the best performer on ProofWriter, SIPDO achieves highly competitive results on ProofWriter, trailing by only 0.4% on GPT-4o-mini and 2% on GPT-4o, underscoring its strong adaptability to structured reasoning tasks. Crucially, unlike neuro-symbolic approaches that rely on predefined rule-based datasets, SIPDO is trained entirely on generated synthetic data, demonstrating the effectiveness of LLM-driven data augmentation for enhancing logical inference across diverse reasoning benchmarks. SIPDO outputs a revised prompt that surpasses baselines, validating its effectiveness for prompt design and performance gains. All generated prompts appear in Appendix B.
@@ 4.3 Results and Analysis
Tasks Vanilla Neuro-S CoT REVOLVE SIPDO Baseline Neuro-S CoT REVOLVE SIPDO
@@
FOLIO 71.2 79.2 72.6 65.7 83.2($\uparrow 4.0$) 51.2 73.2 69.3 62.8 81.1($\uparrow 7.9$)
PrOntoQA 80.4 85.2 95.6 85.4 96.3($\uparrow 0.7$) 74.6 79.3 89.3 83.4 91.3($\uparrow 2.0$)
Average 70.0 82.0 80.2 68.4 86.4($\uparrow 4.2$) 59.5 77.4 73.5 64.9 83.9($\uparrow 6.5$)
Table 3: Performance(%) on ProofWriter, FOLIO, and PrOntoQA by Neuro-Symbolic, CoT, REVOLVE, SIPDO, and Baseline Prompting methods across GPT-4o and GPT-4o-mini.
Model Penguins Geome. Epistemic Obj.Cnt. Temporal Causal Avg.
@@
4.4 Ablation Study
Difficulty Gradient. To assess the contribution of the difficulty gradient, we conduct an ablation study by comparing without difficulty level. As Table 4 shows, every BIG-Bench sub-task suffers when the difficulty gradient is absent. On average, GPT-4o loses $17.3\%$ accuracy, while the weaker GPT-4o-mini drops $24.3\%$, confirming that smaller models depends even more on the difficulty gradient. The steepest declines appear on tasks Object Counting (40.9 % and 54.4 %) and Geometric Shapes (17.2 % and 35.1 %). Even comparatively simple tasks—Temporal Sequences and Epistemic Reasoning—still lose up to 6 %. These results indicate that within the OpenAI model family, the weaker model is more sensitive to the absence of a difficulty gradient and the benefit of a progressive difficulty schedule becomes more pronounced. Without this gradient, generated prompts tend to be shorter and easier, often failing to capture complex reasoning patterns (details in Appendix C).
One-Shot Extremes. We experimented with replacing the step-wise difficulty gradient by a one-shot extremes sampler that tells the generator to create the most unusual examples. On our synthetic suites this shortcut delivered no measurable gain. The “extreme” samples were either solved instantly or only slight perturbations of original cases, leaving the optimizer with no fresh errors to exploit. We suspect the idea will pay off in real-world corpora (e.g. financial statements) where genuine edge cases abound and can expose blind spots that our synthetic tasks do not capture.
@@ 5 Conclusion
We presented SIPDO, a data-centric framework that converts augmentation into a live feedback signal for prompt optimization. A generator creates progressively harder examples, and an auto-prompt optimizer uses them to expose and correct prompt weaknesses. This coupling produces consistent accuracy gains across diverse reasoning benchmarks, outperforming several leading baselines. Beyond these empirical results, SIPDO shows how data-generation strategies and prompt optimization can reinforce one another, linking ideas from curriculum learning and adaptive optimization with LLM practice. Further investigation on domain specific corpora such as financial filings and clinical notes and exploration of fully automated variants that refine prompts through continuous model feedback will clarify SIPDO’s broader value and extend its principles to new settings.
Appendix A Detailed Ablation Table Studies
In this section, we provide more comprehensive and detailed ablation studies that demonstrate our studies and findings.
Task Difficulty GPT-4o-mini GPT-4o Time (s) Money (\$) Tokens Acc. (%) Time (s) Money (\$) Tokens Acc. (%) Penguins 5 217.43 0.0130 67269 83.7 58.55 0.1089 35426 91.3 10 1078.47 0.0929 527358 92.59 561.05 1.2100 410807 93.33 15 2226.68 0.2174 1250646 92.59 2392.23 6.4528 2275582 96.0 20 19753.99 2.6698 16695611 93.33 3452.01 7.8194 2759306 96.67 Epistemic 5 49.07 0.0027 14297 75.33 119.28 0.1401 42985 78.0 10 686.18 0.0420 227607 76.0 555.86 0.6814 224480 79.0 15 1471.69 0.1149 640713 80.33 1440.52 2.0357 695678 83.67 20 3279.38 0.2583 1483605 80.0 2885.80 4.0457 1418053 84.0 Geometric 5 51.70 0.00369 19079 35.33 106.06 0.0955 27548 50.0 10 690.22 0.03237 144212 49.17 560.95 0.4726 133780 67.67 15 1326.86 0.06710 296540 72.22 964.28 0.9158 268912 80.0 20 2503.59 0.20610 571178 78.06 1334.71 1.4882 450575 87.78 causal_judgement 5 103.80 0.0064 35013 58.42 149.56 0.1611 52383 67.95 10 899.33 0.0699 385346 66.66 156.22 0.2761 101715 81.6 15 2066.64 0.2017 1152961 79.67 534.45 0.6921 237519 82.42 20 443.71 0.0388 236561 67.89 3033.81 4.8447 1738332 79.49 temporal 5 35.40 0.0016 7650 94.0 39.27 0.0283 7813 99.67 10 147.00 0.0066 26538 97.5 647.34 0.8005 267894 99.33 15 2166.98 0.1876 1039249 99.67 1432.08 2.1206 736935 99.33 20 1518.23 0.1187 630972 99.64 2235.50 3.5084 1255135 99.66 object counting 5 190.52 0.0122 60765 91.0 115.30 0.1660 49776 97.0 10 1094.08 0.0684 363795 92.0 512.40 0.7296 232588 98.0 15 2592.44 0.2069 1148406 96.0 1396.33 1.8831 624918 99.33 20 5078.75 0.3968 2260081 96.0 2848.81 4.3167 1459739 99.33
Table 5: SIPDO’s cost (time, money, tokens) and accuracy across tasks and difficulty levels for GPT-4o-mini and GPT-4o.
Model Method Penguins Epistemic Geometric Temporal Causal Judgment Object Counting Run Time (s) Acc. (%) Run Time (s) Acc. (%) Run Time (s) Acc. (%) Run Time (s) Acc. (%) Run Time (s) Acc. (%) Run Time (s) Acc. (%) GPT-4o PromptAgent 18300 96.1 23675 91.6 30639 83.0 40399 98.4 18232 77.8 37453 88.2 SIPDO 2392 96.4 556 86.3 1335 82.2 647 99.3 156 79.0 1094 91.1 GPT-4o-mini PromptAgent 9398 89.8 24150 86.0 46012 72.0 48561 94.6 15097 84.6 41977 84.3 SIPDO 2227 92.1 686 85.1 2504 73.2 147 98.0 899 88.0 512 87.5
Table 6: Run time and accuracy across tasks for SIPDO and PromptAgent
Cost in time (s) Cost in money (\$) Performance (%) Task Remove voters With voters Remove voters With voters Remove voters With voters College Biology 763 912 0.0167 0.0221 95.14 96.5 Machine Learning 375 680 0.0096 0.7974 76.79 93.8 College Computer Science 432 1311 0.0104 1.8080 88.00 93.0
Table 7: Impact of SIPDO’s voters on cost and performance across tasks.
Accuracy (%) Cost in time (s) Task w/o generator w/o error analysis w/o revisor w/o generator w/o error analysis w/o revisor College Biology 95.83 95.83 95.14 331 677 314 Machine Learning 90.18 84.82 77.68 767 1622 353 College Computer Science 94.00 90.00 89.00 2405 2852 2861
Table 8: Ablation results on accuracy and time cost when removing each component of SIPDO.
Appendix B Optimzed Prompts for different tasks